# UPose3D: Uncertainty-Aware 3D Human Pose Estimation with Cross-View and Temporal Cues

### v2

@@ IV-B Evaluation Metrics
We adopt the standard evaluation metrics used in prior works for our InD experiments on the Human3.6m [11] dataset. Mean Per Joint Position Error (MPJPE) represents the distance between the ground truth and predicted pose in millimeters, while its Procrustes Aligned version (PA-MPJPE) and the Normalized variant (N-MPJPE) show how well the predicted pose fits the ground-truth keypoints after a similarity transformation. Next, we report the translation aligned error (TA-MPJPE) and PA-MPJPE for our OoD experiments on the RICH [12] dataset following prior works [49, 50, 51, 27, 52, 53].
IV-C Implementation Details
We implement our pipeline using PyTorch [54] and MMPose [55]. To compare our method with prior works, we choose a CPN [36] backbone based on ResNet152 [56] with an input size 384 $\times$ 384 and an RLE head [23] for 2D pose estimation. We fine-tune our 2D pose estimator jointly on Human3.6m [11] and MPII [57] datasets starting from the weights of the checkpoints provided in [2]. For our OoD experiments, we use an HRNet-W48 [24] with 384 $\times$ 384 input size and train it from scratch.
@@ IV-C Implementation Details
In our point cloud encoder, we use 4 multi-head self-attention blocks, each with 4 attention heads. Additionally, we reduce its hidden dimension to 64 to accommodate the relatively smaller size of our point clouds. Following the point cloud encoder, the features are concatenated with two positional embeddings of size 64 to represent the time frames and different skeleton joints. Our spatiotemporal encoder employs four criss-cross transformer blocks [33, 8] with 4 attention heads each. We set the hidden layer dimension of our transformers to 64 and the hidden dimension of prediction heads to 1024. We then train our pose compiler module using AdamW optimizer [58] with a batch size of 64 on the AMASS [13] dataset for both OoD and InD evaluations. We use a learning rate of 4e-5 with a warm-up factor of 1e-4 for the first 500 iterations and a cosine annealing scheduler over the next 20,000 iterations. The network training takes about 6 hours on an NVIDIA 2080 RTX GPU.
@@
TABLE I: The comparison of our method in InD settings against prior multi-view works on the full test set of the Human3.6m dataset. (-) denotes that the error was not reported in the original work. The errors are reported in mm.
@@
TABLE II: Comparison of our method in OoD setting on the Rich [12] dataset against prior works. * denotes our implementation of prior works.
@@
AdaFuse* [4] 524.0 85.8
HRNet-W48+RANSAC* 338.9 138.9
HRNet-W48+DLT* 66.0 55.1
Ours ($T=1$) 36.2 33.4
Ours ($T=27$) 34.7 32.0
IV-D Baselines
On Human3.6m [11], we compare our UPose3D with methods that infer 3D human keypoints using only 2D supervised models [36, 3, 7, 1, 63, 17]. Additionally, we provide a summary of methods that rely on 3D annotations [62, 61, 6, 32, 4, 60, 3, 2, 7] that use ground-truth camera parameters during inference, except for a few [62, 61, 10] that use uncalibrated cameras. We also report a more in-depth comparison with weakly-supervised and semi-supervised approaches [65, 17, 66, 59, 67], camera-based multi-view methods that do not rely on camera parameters [62, 61, 10], and monocular 3D pose estimation methods [53, 52, 27, 51, 50, 49]. We also report and compare the performance of our method against triangulation approaches, such as DLT, RANSAC, and RPSM [7]. On the RICH dataset [12], we compare our model with triangulation approaches and replicate the performance of AdaFuse [4], one of the top-performing methods of Human3.6m dataset [11].
@@ IV-D Baselines
TABLE III: Additional comparisons with prior works on the full test set of the Human3.6m dataset in InD settings. (-) denotes that the error was not reported in the original work.
@@
TABLE IV: Additional comparison of our method in OoD setting on Rich dataset against prior works. * denotes our replication of prior works.
@@
SkelFormer [46] 44.2 35.6 SMPL
@@
V Results
V-A In-Distribution Performance
In Table I, we present a comparison of our InD results with other state-of-the-art methods on the Human3.6m dataset, divided into two categories: methods that are trained directly with 3D annotated data and methods like ours that leverage only 2D supervision. We observe that our method outperforms other methods that rely on 2D supervision. In addition, UPose3D yields results that are competitive with state-of-the-art methods relying on 3D supervision. For example, we achieve similar results to MTF-Transformers, which, like our pose compiler, takes in multi-view and temporal 2D keypoints as input, but estimates the 3D poses directly using a deep network with 10.1M parameters. This indicates that our proposed pose compiler module and training strategies are effective in obtaining precise keypoint position estimations.
We also experiment with different 2D pose estimator backbones. We see that the choice of backbone is an important factor for accuracy and that CPN is preferred over ResNet152. Interestingly, a vanilla triangulation algorithm (DLT) can outperform prior works given accurate 2D predictions from strong pose estimators. Combining strong pose estimators with our method yields even better performance. For instance, UPose3D improves the MPJPE by 3.6 mm when compared to vanilla triangulation using a CPN backbone (DLT+CPN). Finally, we observe that increasing the temporal frame window positively impacts performance, reducing MPJPE by 0.5 mm when using 27 frames instead of a single frame.
V-B Out-of-Distribution Generalization
In order to evaluate the performance of UPose3D in OoD settings, we compare it to the best-performing baseline on the Human3.6m [11] dataset, namely AdaFuse [4]. Additionally, we report the results for triangulation techniques applied to 2D pose estimators Table II. In this experiment setup, none of the models nor their components are trained on the RICH [12] dataset. First, we observe that our method obtains the best results, achieving half of the error of the next best method. Additionally, we observe that AdaFuse performs poorly on this dataset despite being considered state-of-the-art on Human3.6m [11] dataset. Our analysis of AdaFuse performance on different clips of RICH [12] dataset indicates that this method performs well only if all the viewpoints’ predictions are within a reasonable range, whereas occasional noisy predictions cause large triangulation errors. In conclusion, we believe that by training on limited in-studio datasets, methods such as AdaFuse fail to generalize well to outdoor and in-the-wild environments. In contrast, our approach uses synthetic data and performs consistently when applied to unseen datasets.
V-C Additional Baselines and Comparisons
Table III complements Table I by providing more comparisons with prior works on the Human3.6m [11] dataset. Here, we include 3D keypoint estimation approaches regardless of their input modality or supervision type in InD settings. We observe that our method outperforms all of the other approaches despite only using 2D supervision. Additionally, in Table IV, we compare our work with prior research on the RICH [12] dataset. Since this dataset was recently published, only monocular 3D body modeling techniques have reported their performance on this dataset. Here, we observe that our method outperforms the majority of prior works. More importantly, when comparing Table III and Table IV we notice that our method performs consistently on InD and OoD data, showing generalizability across in-studio and outdoor environments.
TABLE V: Ablation experiments on the Human3.6m dataset with $T=27$.
[…]
UPose3D 26.42 23.42
     w/o compiler 37.14 33.90
     w/o image branch 69.90 50.97
      w/o compiler uncertainty 26.42 23.58
      w/o image uncertainty 27.61 24.88
      w/o uncertainty 48.11 41.20
         w/o image branch 77.25 54.02
         w/o compiler 30.48 27.63
V-D Ablation Study
To investigate the impact of each component of UPose3D, we systematically remove them and report the results of our method with a CPN backbone on the test set of Human3.6m [11]. We perform all our experiments using a model trained with a temporal window of 27 frames. Accordingly, we first ablate the effect of our pose compiler and uncertainty modeling. Next, we study the details of our pipeline to evaluate its performance under different inputs, network architectures, and initialization strategies for 3D keypoint estimation. We further investigate the impact of temporal length, our spatiotemporal encoder’s architecture, different formulations of the point clouds, and our initialization strategy for 3D keypoint estimation.
TABLE VI: Additional ablation study on Human3.6m dataset. We only report the computation cost of our pose compiler and exclude the CPN [36] network (with 5.16T FLOPs and 64.87M parameters) for 27 frames of 4 views.
[…]
     w/ zero init 28.51 32.85 65.407 12.5 2.04
[…]
     w/ T = 243 33.17 25.11 65.430 10.9 20.18
     w/ concurent attention 26.57 23.61 65.391 10.3 2.01
     w/ full attention 26.50 23.57 65.322 10.3 2.28
     w/ full attention (T = 243) 34.97 28.60 65.336 10.3 51.56
     w/ epipolar line 26.46 23.45 65.407 10.1 2.04
     w/ relative camera pos. emb. 26.37 23.43 65.407 10.2 2.04
Effect of Pose Compiler. Our UPose3D employs the pose compiler module to improve the keypoint and uncertainty predictions using cross-view and spatiotemporal information. By ablating this component, we use the original keypoints and uncertainties predicted by our 2D pose estimator to perform MLE. As shown in Table V, this experiment results in an additional 9.28 $mm$ error resulting in a higher error compared to the DLT algorithm (See Table I). This shows the effectiveness of our pose compiler for 3D pose estimation using normalizing flows.
Effect of Image Branch. We study the impact of the image branch by removing it from the pipeline, which causes a significant rise in the error as seen in Table I. This indicates the significance of keeping the original predictions for the final estimation.
Effect of Uncertainty Modeling. We employ the normalizing flows of our RLE heads during the MLE loss minimization stage to incorporate the uncertainties of our predictions toward improving the model’s robustness. To study the effect of uncertainty modeling, we remove them from the 2D pose estimator and pose compiler branches. Accordingly, we first remove the uncertainties from our pose compiler and replace our likelihood loss function (Eq. 1) with a reprojection distance loss. Table I shows that the PA-MPJPE error slightly rises, but the MPJPE remains unchanged. Similarly in the next experiment, we remove the image pose estimator uncertainties. This ablation experiment results in 1.2 mm rise in the error. Finally, we completely remove the uncertainties from both branches, effectively reducing our problem to a classic triangulation problem without confidence that can be solved via DLT algorithm. As a result, we find the 3D keypoints using a DLT algorithm given the keypoints from each branch. This variation shows over 20 mm higher error than UPose3D. In our final two experiments, we analyze the accuract of each keypoint branch without uncertainty using a DLT triangulation algorithm. The first experiment obtains the highest error in the absence of uncertainties and image pose estimator, while the second experiment results in the same performance of a simple DLT algorithm without cross-view fusion.
Random Initialization. We use L-BFGS [42] optimization algorithm to iteratively solve the 3D keypoint MLE. To speed up this process, we stop the optimization when the changes of our optimization variables, namely $U$, are less than a specific tolerance (Tol.=0.001.=0.001. = 0.001 mm). We speed up the optimization process even further by using a DLT algorithm to initialize the 3D points $U$. Table VI first examines the effect of our initialization strategy when $U$ is initialized to zero, and the tolerance remains unchanged, showing a significant rise in the 3D keypoint estimation error and inference time. Next, Table VI shows that by lowering the tolerance, zero-initialization performs similarly to our proposed strategy, but at 3 times more inference time. Therefore, we conclude that unlike prior works [1], our method is not reliant on initialization and the initialization only speeds up the estimation process. This may be due to the smooth nature of the uncertainty distributions learned by the normalizing flows [23].
Longer Temporal Window. We study the computational cost and performance impact of very long temporal context size. Following [9], we report the performance of UPose3D when 243 frames, as opposed to 27 frames, are used to infer the 3D keypoints of the center frame in Table VI. This new model takes 10 times more FLOPs to compute and does not perform as well as our original model. This may be because our synthetic data augmentations and corruption strategies are tuned for smaller time windows as longer context sizes were not in our considerations. Our observations of the training and validation losses also show signs of overfitting during training for longer time windows. As extremely long context sizes are not in the scope of this paper, we do not tune these models and leave them for future research.
Pose Compiler Architecture. We compare the effect of our criss-cross attention modules with vanilla (full) and concurrent attention. Table VI shows that criss-cross attention outperforms the other two designs while requiring less computation (FLOPs) than full attention. Additionally, we observe that on the extreme case of very long temporal context sizes (T = 243), criss-cross attention still outperforms full attention models by 1.8 mm while requiring 60% less computations.

### v3

@@ 4.2 Evaluation Metrics
We adopt the standard evaluation metrics from prior works for our InD experiments on the Human3.6m [18] dataset. Mean Per Joint Position Error (MPJPE) represents the distance between the ground-truth and predicted pose in millimeters, while its Procrustes Aligned version (PA-MPJPE) and the Normalized variant (N-MPJPE) show how well the predicted pose fits the ground-truth keypoints after a similarity transformation. Next, we report the Translation Aligned error (TA-MPJPE) and PA-MPJPE for our OoD experiments on the RICH [16] dataset following prior works [34, 26, 28, 35, 59, 54].
4.3 Implementation Details
We implement our pipeline using PyTorch [45] and MMPose [6]. Here, we provide the details of our neural networks and the computation costs of our pose compiler. We encourage readers to refer to our Supplementary Materials for information on our multi-view data synthesis and augmentation strategies.
@@ 4.3 Implementation Details
To compare our method with prior works, we choose a CPN [5] backbone based on ResNet152 [13] with an input size 384 $\times$ 384 and an RLE head [31] for 2D pose estimation. We fine-tune our 2D pose estimator jointly on Human3.6m [18] and MPII [1] datasets starting from the weights of the checkpoints provided by [19]. For our OoD experiments, we use an HRNet-W48 [57] with 384 $\times$ 288 input size and train it from scratch. Our point cloud encoder uses 4 multi-head self-attention blocks, each with 4 attention heads. Additionally, we reduce its hidden dimension to 64 to accommodate the relatively smaller size of our point clouds. Following the point cloud encoder, the features are concatenated with two positional embeddings of size 64 to represent the time frames and different skeleton joints. Our spatiotemporal encoder employs four criss-cross transformer blocks [17, 58] with 4 attention heads each. We set the hidden layer dimension of our transformers to 64 and the hidden dimension of prediction heads to 1024. We then train our pose compiler module using AdamW optimizer [40] with a batch size of 64 on the AMASS [42] dataset for both OoD and InD evaluations. We use a learning rate of 4e-5 with a warm-up factor of 1e-4 for the first 500 iterations and a cosine annealing scheduler over the next 20,000 iterations. The network training takes about 6 hours on an NVIDIA 2080 RTX GPU.
@@
4.4 Baselines
On Human3.6m [18], we compare our UPose3D with methods that infer 3D human keypoints using only 2D supervised models [5, 14, 47, 60, 23, 27]. Additionally, we provide a summary of methods that rely on direct 3D annotations [20, 10, 55, 66, 49, 19], except for a few [20, 10, 48] that use uncalibrated cameras. Moreover, we compare our work with methods that only rely on multi-view 2D keypoints along with the camera parameters [67, 41, 14, 47]. Please refer to our Supplementary Materials for a more in-depth comparison with weakly-supervised and semi-supervised approaches [53, 27, 62, 9, 56], multi-view methods without that do not rely on camera parameters [20, 10, 48], and monocular 3D pose estimation methods [54, 59, 35, 28, 26, 34]. We also report and compare the performance of our method against triangulation approaches, such as DLT, RANSAC, and RPSM [47]. On the RICH dataset [16], we compare our model with triangulation approaches and replicate the performance of AdaFuse [66], which is one of the top-performing methods on the Human3.6m dataset [18].
Table 1: The comparison of our method in InD settings against prior multi-view works on the full test set of the Human3.6m dataset. The errors are reported in mm.
@@ 4.4 Baselines
2D Supervision + Camera Parameters
TransFusion [41] ResNet50 1 25.8 - -
DLT* [13] ResNet50 1 71.2 62.8 70.9
UPose3D (Ours)* ResNet50 27 57.3 51.7 57.2
Cross-view Fusion [47] ResNet152 1 26.2 - -
Epipolar Transformers [14] ResNet152 1 19.0 - -
TR Loss [67] HRNet [57] 1 25.8 - -
UPose3D (Ours) CPN 27 26.2 23.1 25.4
- denotes the error was not reported in the original work.
* denotes training from scratch with no additional data.
✝ denotes using 1% of unlabeled testing data during training.
5 Results
5.1 In-Distribution Performance
In Tab. 1, we present a comparison of our InD results with other state-of-the-art methods on the Human3.6m dataset, divided into three categories: a) methods that, unlike ours, are trained directly with 3D annotated data; b) methods that leverage only 2D supervision; and c) methods that rely on 2D supervision along with the camera parameters. We observe that UPose3D outperforms all other methods that rely on 2D supervision. In addition, it yields results that are competitive to state-of-the-art methods relying on 3D supervision. For example, our method achieves similar results to MTF-Transformers [55], which, like our pose compiler, takes in multi-view and temporal 2D keypoints as input but estimates the 3D poses directly using a deep network with 10.1M parameters. Furthermore, among methods that use the camera parameters of the Human3.6m dataset, our method performs similarly to TR Loss [67], which uses 1% of the unlabeled testing set during training. Although Epipolar Transformers [14] perform better than ours due to their effective feature fusion strategy, it relies on precise camera parameters and has constraints on the viewing angle of neighbouring cameras. The consistent performance of our approach in both groups indicates the effectiveness of our proposed pose compiler and training strategies in obtaining precise keypoints regardless of the target dataset’s camera setup.
We also show that the choice of the 2D keypoint estimator backbone is an important factor for accuracy and that CPN is preferred over ResNet152. Interestingly, a vanilla triangulation algorithm, DLT, can outperform prior works given accurate 2D predictions from strong pose estimators. Combining strong pose estimators with our method yields even better performance. For instance, UPose3D improves the MPJPE by 3.6 mm compared to vanilla triangulation using a CPN backbone (DLT+CPN). Additionally, we experiment with a ResNet-50 model trained from scratch in the third group (2D supervised with camera parameters) and observe a 19.5% improvement when UPose3D is used. Finally, we demonstrate the positive impact of the temporal frame window, which reduces the MPJPE by 0.5 mm when using 27 frames instead of a single frame.
5.2 Out-of-Distribution Generalization
To evaluate the performance of UPose3D in OoD settings, we compare it to the best-performing baseline on the Human3.6m [18] dataset, namely AdaFuse [66]. Additionally, we report the results for triangulation techniques applied to 2D pose estimators in Tab. 2. In this experiment setup, neither the models nor their components are trained on the RICH [16] dataset. First, we observe that our method, despite the training source dataset, obtains the best results by achieving half of the error of the next best method. Additionally, we observe that AdaFuse performs poorly on this dataset despite being considered state-of-the-art on the Human3.6m [18] dataset. Our analysis of AdaFuse performance on different clips of RICH [16] indicates that this method performs well only if all the viewpoints’ predictions are within a reasonable range, whereas occasional noisy predictions cause large triangulation errors. In contrast, our solution deduces the correct pose by effectively discarding views with low confidence. In conclusion, when training on in-studio datasets, methods such as AdaFuse fail to adapt to in-the-wild environments. At the same time, our approach uses uncertainty modeling to better generalize to unseen camera configurations. UPose3D is not limited to the source dataset skeleton configuration because synthetic data are used for training, and it performs consistently between InD and OoD evaluations.
Table 2: Comparison of our method in OoD setting on the RICH [16] dataset against prior works. The training sources are denoted as (2D pose estimator, Pose compiler).
@@ 5.2 Out-of-Distribution Generalization
AdaFuse* [66] (Human3.6m+MPII, Human3.6m) 524.0 85.8
Ours ($T=27$) (Human3.6m+MPII, Human3.6m) 51.8 43.6
Ours ($T=1$) (Human3.6m+MPII, AMASS) 49.9 42.3
HRNet-W48+DLT* (COCO, N/A) 66.0 55.1
HRNet-W48+Grid Search* (COCO, N/A) 64.4 54.9
Ours ($T=1$) (COCO, AMASS) 36.2 33.4
Ours ($T=27$) (COCO, AMASS) 34.7 32.0
* denotes our implementation of prior works.
5.3 Ablation Study
To investigate the impact of each component, we systematically remove them during inference and report the results with a CPN backbone on the test set of Human3.6m [18]. We perform all experiments using a model with a temporal window of 27 frames. Please refer to our Supplementary Materials for additional analysis on temporal length, spatiotemporal encoder’s architecture, different point cloud formulations, and initialization strategies for 3D estimation.
Pose Compiler. We employ the pose compiler module to improve the keypoint and uncertainty predictions using cross-view and spatiotemporal information. By ablating this component, we use the original keypoints and uncertainties predicted by our 2D pose estimator to perform MLE. As shown in Tab. 3, this experiment results in an additional 9.28 mm error resulting in a higher error compared to the DLT algorithm (see Tab. 1). This shows the effectiveness of our pose compiler for 3D pose estimation using normalizing flows.
Table 3: Ablation experiments on the Human3.6m dataset with $T=27$.
[…]
UPose3D 26.42 23.42
      w/o compiler 37.14 33.90
      w/o image branch 69.90 50.97
       w/o compiler uncertainty 26.42 23.58
       w/o image uncertainty 27.61 24.88
       w/o uncertainty 48.11 41.20
          w/o image branch 77.25 54.02
          w/o compiler 30.48 27.63
Image Branch. Removing the image branch from the pipeline causes a significant rise in the error, seen in Tab. 1, indicating the significance of keeping the original predictions for the final estimation.
Uncertainty Modeling. We employ the normalizing flows of our RLE heads during the MLE loss minimization stage to incorporate the uncertainties of our predictions toward improving the model’s robustness. We employ the RLE head’s normalizing flows module to model the uncertainties, improving robustness. To study the effect of uncertainty modeling, we remove the RLE heads from the 2D pose estimator and pose compiler branches. Accordingly, we first remove the uncertainties from our pose compiler and replace our likelihood loss function (Eq. 1) with a reprojection distance loss. Table 1 shows that the PA-MPJPE error slightly rises, but the MPJPE remains unchanged. However, it should be noted that removing the pose compiler uncertainty during training causes more performance degradation by resulting in 26.8 mm, which shows the significance of uncertainty modeling. Next, we remove the image pose estimator uncertainties, which results in a 1.2 mm rise in the error. Finally, we completely remove the uncertainties from both branches, effectively reducing our problem to a classic triangulation problem without confidence that can be solved via the DLT algorithm. This variation shows over 20 mm higher error than UPose3D. In our final two experiments, we analyze the image and pose compiler predictions without uncertainty modeling. The first experiment shows the highest error, while the second experiment results in the same performance as a simple DLT algorithm without cross-view fusion.
5.4 Computation Costs
In Tab. 4, we compare the total computation costs of our method in comparison to state-of-the-art 3D approaches in single-frame scenarios. Compared to others, UPose3D has fewer parameters and significantly less computational cost, especially with more cameras. The computational cost of our 10-frame model can be broken down to 508.5G FLOPs for CPN, 0.385G for pose compiler, and 8.8$\pm$1.25G for the optimization process. Due to the computational cost fluctuations caused by the L-BFGS optimizer, we measure it by averaging 20 runs using randomly selected cameras from the HUMBI dataset [64, 65].
Table 4: Comparison of computation costs.
Method Params(M)$\downarrow$ FLOPs for 4 Cams(G)$\downarrow$ FLOPs for 10 Cams(G)$\downarrow$
Learnable Triangulation [19] 80.6 716.1 1326.9
Epipolar Transformers [14] 68.1 406.5 1016.2
MTF-Transformers [55] 78.6 407.0 1017.8
AdaFuse [66] 69.7 595.0 1487.6
Ours 65.4 208.7 517.7
Figure 3: We demonstrate the scalability of UPose3D to the number of cameras.
Figure 4: Illustration of UPose3D on Human3.6m [18] (right) and RICH [16] (left) datasets, showing the accurate 3D pose estimated by our UPose3D (top) compared to ground-truth (bottom).
5.5 Viewpoint Scalability
In the context of multi-view 3D human pose estimation, scalability may be defined as the capability of a method to handle data from an increasing number of camera views effectively. Specifically, such a solution should have an almost constant runtime with constantly decreasing error when more cameras are used. Accordingly, to test the scalability of our method, we evaluate it with different numbers of cameras on the CMU Panoptic [22] and the HUMBI [64, 65] datasets and present the results in Fig. 3. For this experiment, we use the HRNet-W48 [57] backbone, similar to our experiments on the RICH dataset, to analyze scalability in OoD scenarios. Additionally, we report and compare our error and runtime with two triangulation techniques, RANSAC and DLT. We tune the RANSAC implementation of [19] and run it for 20 iterations. The error analysis shows that as the number of views increases, UPose3D’s error continuously improves while DLT and RANSAC reach a plateau. Next, we discard the common factor of the 2D pose estimators for runtime scalability analysis. The inference time of our pipeline with a single batch size remains constant despite the number of cameras. However, increasing the batch size causes runtime improvements at the cost of memory and runtime variability.
5.6 Qualitative Analysis
Fig. 4 shows the outputs of our method on challenging examples, where we sample the keypoint distributions in every pixel to demonstrate the keypoint likelihoods from the 2D pose estimator and our pose compiler as heatmaps. Despite choosing challenging scenes, we observe that our predictions are still close to the ground-truth keypoints, indicating that our method does not produce any significant outliers within its output even on the unseen samples of RICH [16].
6 Conclusion
This paper presents UPose3D, a multi-view 3D human pose estimation method designed to address the challenges in generalization, scalability, and over-reliance on real-world 3D annotated data. It includes a novel cross-view fusion strategy that scales well with varying camera numbers and volume sizes. Additionally, UPose3D integrates a pose compiler that learns to predict keypoint positions and uncertainties given the cross-view and temporal information. Finally, the uncertainties and predicted key points from two sources of the image branch and pose compiler are used for uncertainty-aware 3D pose estimation. UPose3D outperforms state-of-the-art approaches in OoD setup and achieves competitive InD results without any 3D pose annotations. Therefore, UPose3D may positively impact applications where performance and robustness are crucial.
Future Work. Our primary focus in this paper was the fidelity of estimated output 3D poses from multi-view inputs without requiring 3D supervision. We acknowledge, however, that our method could be optimized for real-time applications. A promising avenue to achieve faster inference involves the exploration of specialized second-order optimizers that use a deep-learning neural network to estimate the Hessian matrix during likelihood maximization. Furthermore, depth and trajectory modalities can be explored for additional supervision in the likelihood function to reduce noisy predictions.
Acknowledgements
This work was performed during Vandad Davoodnia’s internship at Ubisoft Laforge, partially funded by Mitacs through the Accelerate program.
Appendix 0.A Additional Details
0.A.1 Details on Training and Multi-view Data Synthesis
Let $r\in\mathbb{R}^{T\times 3}$, $\Theta\in\mathbb{R}^{T\times 55\times 3\times 3}$, and $\beta\in\mathbb{R}^{16}$ be root position, body joint rotation matrices in a temporal window of length $T$, and shape parameters from human motion capture data. We begin our multi-view data generation by augmenting the shape parameters with Gaussian noise with a standard deviation equal to the standard deviation of all shape parameters within the datasets. Next, we apply mediolateral mirroring with a 50% chance and randomly rotate the motion sequence around its center. We pass the augmented $\{r,\Theta,\beta\}$ parameters to the forward-kinematic layer of the SMPL body model to obtain 3D vertices. Lastly, we use a dataset-specific joint regressor on the vertices to extract the 3D keypoints used in the next steps of our pipeline.
After obtaining the camera intrinsic and extrinsic parameters, we project the 3D body keypoints onto each camera view (see Fig. 5-b). We use these 2D keypoints as ground truths $\mu_{g}$ to train our pose compiler (see Fig. 5-c). We then add 2D point corruptions to the 2D keypoints, including Gaussian noise with varying standard deviations, simulated occlusions with varying sizes and probabilities, mediolateral flipping, and occasional truncation effects (see Fig. 5-d). Next, we obtain the cross-view projected keypoints (see Fig. 5-e) via the algorithm described in Sec. 3.2 in the main paper. Finally, we train our pose compiler using the ground-truth keypoints and point clouds containing noisy 2D data, as depicted in Fig. 6.
Figure 6: We illustrate the training routine of our pose compiler using synthetic data generated based on real motion capture sequences.
0.A.2 Details on Criss-cross Attention
As discussed in Sec. 3.3 in the main paper, we use criss-cross attention blocks in our spatiotemporal encoder to process information more efficiently. Accordingly, the cross-view input features $\{f_{i}\}$ are first projected into queries, keys, and values ($Q,K,V\in\mathbb{R}^{T\times J\times 2H}$) via a linear layer. Next, we divide them into temporal $Q_{T},K_{T},V_{T}\in\mathbb{R}^{T\times J\times H}$ and spatial groups $Q_{S},K_{S},V_{S}\in\mathbb{R}^{T\times J\times H}$. The temporal and spatial (skeleton joints) attentions are then calculated in two separate self-attention modules and concatenated before the next feed-forward layer and normalization. As a result of this operation, the receptive field of each transformer layer is the information residing on the spatial and temporal axis, and stacking multiple layers can approximate the full spatiotemporal attention without large computational overhead. In the following sections, we study the effectiveness of our design choice and compare its computation cost and performance against full attention and concurrent attention designs.
Appendix 0.B Additional Experiments and Results
This section describes the 2D datasets used during training and fine-tuning of our 2D pose estimator. We then study the details of our pipeline to evaluate its performance under different inputs, network architectures, and initialization strategies for 3D keypoint estimation. Next, we provide additional comparisons on the Human3.6m [18] dataset with weak or semi-supervised methods. We will also provide more comparisons with monocular pose estimation approaches on the RICH [16] dataset.
Table 5: Additional ablation study on Human3.6m dataset. We only report the computation cost of our pose compiler (in FLOPs) and exclude the CPN [5] network with 5.16T FLOPs for 27 frames of 4 views. Additionally, 64.87M of parameters in all experiments belong to the CPN network.
[…]
     w/ zero init 28.51 32.85 65.407 12.5 2.04
[…]
     w/ T = 243 33.17 25.11 62.660 10.9 20.18
     w/ concurent attention 26.57 23.61 65.391 10.3 2.01
     w/ full attention 26.50 23.57 65.322 10.3 2.28
     w/ full attention (T = 243) 34.97 28.60 65.336 10.3 51.56
     w/ epipolar line 26.46 23.45 65.407 10.1 2.04
     w/ relative camera pos. emb. 26.37 23.43 65.407 10.2 2.04
0.B.1 2D Datasets
COCO WholeBody. The COCO WholeBody [21] dataset is a large-scale whole-body pose estimation dataset with over 250K samples. This dataset is an extension of Common Objects in COntext (COCO) [36] dataset with the same training and testing splits. The dataset provides 133 2D keypoints (17 for body, 6 for feet, 68 for face, and 42 for hands) on in-the-wild images. We use this dataset to train our 2D pose estimator during OoD experiments.
MPII. The MPII Human Pose dataset [1] dataset is a popular 2D pose estimation benchmark. It contains over 40,000 images of people performing over 400 actions in diverse scenarios. The dataset contains 16 body joint labels and is frequently used to pre-train [66] or improve cross-dataset generalization[69, 47]. We use this dataset for 2D pose estimator pre-training and fine-tuning.
0.B.2 Additional Ablation Study
Following the ablation study originally presented in Sec. 5.3 of the main paper, we investigate the impact of temporal length, our spatiotemporal encoder’s architecture, different formulations of the point clouds, and our initialization strategy for 3D keypoint estimation in Tab. 5. Additionally, we report and provide the computational cost comparisons for a single input batch with T = 27 and 4 views. Our pose compiler is significantly smaller than a single 2D pose estimator, taking less than 1% of the total parameter count.
Random Initialization. We use the L-BFGS [37] optimization algorithm to solve the 3D keypoint MLE iteratively. To speed up this process, we stop the optimization when the changes of our optimization variables, namely $U$, are less than a specific tolerance (Tol.=0.001.=0.001. = 0.001 mm). We further speed up the optimization process by using a DLT algorithm to initialize the 3D points $U$. Table 5 first examines the effect of our initialization strategy when $U$ is initialized to zero, and the tolerance remains unchanged, showing a significant rise in the 3D keypoint estimation error and inference time. Next, Tab. 5 shows that by lowering the tolerance, zero-initialization performs similarly to our proposed strategy, but at 3 times more inference time. Therefore, we conclude that unlike prior works [60], our method is not reliant on initialization, and the initialization only speeds up the estimation process. This may be due to the smooth nature of the uncertainty distributions learned by the normalizing flows [31].
Longer Temporal Window. We study the computational cost and performance impact of very long temporal context size. Following [70], we report the performance of UPose3D when 243 frames, as opposed to 27 frames, to infer the 3D keypoints of the center frame in Tab. 5. This new model takes 10 times more FLOPs to compute and does not perform as well as our original model. This may be because our synthetic data augmentations and corruption strategies are tuned for smaller time windows, as longer context sizes were not in our considerations. Our observations of the training and validation losses also show signs of overfitting during training for longer time windows. As extremely long context sizes are not in the scope of this paper, we do not perform any additional tuning of these models and leave them for future research.
Pose Compiler Architecture. We compare the effect of our criss-cross attention modules with vanilla (full) and concurrent attention. Table 5 shows that criss-cross attention outperforms the other two designs while requiring less computation (FLOPs) than full attention. Additionally, we observe that on the extreme case of very long temporal context sizes (T = 243), criss-cross attention still outperforms full attention models by 1.8 mm while requiring 60% less computations.
@@ 0.B.2 Additional Ablation Study
Table 6: Additional comparisons with prior works on the full test set of the Human3.6m dataset in InD settings. (-) denotes that the error was not reported in the original work.
@@
Table 7: Comparison of our method in OoD setting on RICH dataset against prior works. * denotes our replication of prior works.
@@
AdaFuse* [66] 524.0 85.8 3D Keypoints
HRNet-W48+Grid Search* 64.4 54.9 3D Keypoints
HRNet-W48+DLT* 66.0 55.1 3D Keypoints